OpenStudy (anonymous):

is the sequence geometric? If so, identify the common ratio. one-third, two-ninths, 4 over 27, 8 over 81, 16 over 243, ... (1 point) yes; two over three yes; one ninth yes; one sixth not geometric

OpenStudy (texaschic101):

(2/9) / (1/3) = 2/9 * 3 = 6/9 = 2/3 (4/27) / (2/9) = 4/27 * 9/2 = 36/54 = 2/3 (8/81) / (4/27) = 8/81 * 27/4 = 216/324 = 2/3 (16/243) / (8/81) = 16/243 * 81/8 = 1296/1944 = 2/3 Yes....each term is multiplied by 2/3 to get the next term